The sum of two numbers is 66. The larger number is three more than twice the smaller number. Find the numbers.
Answer by checkley79(3341)

X+(2X+3)=66
3X=66+3
3X=63
X=63/3
X=21 ANS. FOR THE SMALLER NUMBER.
2*21+3=42+3=45 ANS. FOR THE LARGER NUMBER.
PROOF:
21+45=66
66=66